OpenStudy (lgbasallote):

Prove: \[\huge \frac{2\tan x}{1 + \tan^2 z} = \sin (2x)\]

OpenStudy (phi):

you could memorize 1+tan^2 identity or you could change tan to sin/cos and 1 to cos^2/cos^2 and add, simplify use the double angle formula or sin(a+b)= sin(a)cos(b)+ sin(b)cos(a) for a=b this becomes 2 sin(a)cos(a)

OpenStudy (anonymous):

\[Sin(2x) = 2sinxcosx\]

OpenStudy (anonymous):

\[2sinxcosx=\frac{2sinxcosx }{ \cos^2x+\sin^2x }\]

OpenStudy (anonymous):

nw Divide Numerator and numerator by cos^2x

OpenStudy (anonymous):

\[=\frac{ 2tanx }{ 1+\tan^2x }\] hence proved...)

OpenStudy (lgbasallote):

that's x not z...darn this lag

OpenStudy (lgbasallote):

@phi is it possible to substitute sec^2 x?

OpenStudy (phi):

yes, there is more than one way to do this